I need help with solving systems of equations
So the first method i need help with is elimination, here is the problem
{15x-5y=30
{y=2x+3
Then i need help with the second method the graphing method
it is the same system as you see above

Answer by ewatrrr(24785)

 
Hi,
doing the Graphing First
y = 2x+3 green p(0,3) and P(3,9) Plot, connect with line
y = 3x-6 blue P(0, -6) and P(2, 0) Plot, connect with line
Line intersect at(9,21), x = 9, y = 21 the solution

Let's see
15x - 5y = 30
2x - y = -3 Multiply by 5 and subtracting from 1st EQ to eliminate y
5x = 45
x = 9 and using y = 2x+3, y = 21